Question 3 3 pts Consider a company with the following reported net income (i.e. reported sales) and accounts receivable. Compute the cash flows in each of the three years and report the NPV of these cash flows. Use a 7.5% discount rate. Assume Accounts Receivable in year 0 is $0. Year Net Income Accounts Receivable 2 $4 $10 3 $4 $3 $3 $2. Question 4 3 pts A firm has reported deferred tax liabilities of $193 and $294 this year and last year, respectively. Its GAAP after-tax cash flow was $1,000 this year. What is its actual after-tax cash flow this year?Please answer both

Carry out calculations to at least 4 decimal places. Enter percentages as whole numbers. Example: 3.03% should be entered as 3.03. Do not include commas or dollar signs in numerical answers.
Answer #1

3)

cash flow from operations = net income + changes in account receivables

cash flow in year 1 = CF1 = net income + (account receivables in year 1 - account receivables in year 0) = 3 + (2-0) = 5

cash flow in year 2 = CF2 = net income + (account receivables in year 2 - account receivables in year 1) =  4 + (10-2) = 12

cash flow in year 3 = CF3 = net income + (account receivables in year 3 - account receivables in year 2) =  4 + (3-10) = -3

NPV = CF1/(1+R)1 + CF2/(1+R)2 + CF3/(1+R)3

= 5/(1.075)1 + 12/(1.075)2 - 3/(1.075)3

= 4.651 + 10.384 - 2.414

= 12.621

4)

Differed tax liability is the tax that is due which have to paid in future. Since 193 and 294 are the differred tax liabilities for the last 2 years which have to paid this year, they have to be deducted from this year cash flow to get net cash flow

actual after tax cash flow = 1000-(193+294) = 513